In the ground state of harmonic oscillator,what is the probability of finding the particle outside the classical allowed region
1
Expert's answer
2020-09-28T06:02:18-0400

"P_{classical}=\frac{1}{\sqrt\pi}\intop_{-1}^{1}e^{-\xi^2}d\xi"

This integral is the error function, so we find;

"P_{classical}=erf(1)=0.8427"

The probability of being outside the classical region is then;

"1-P_{classical}=0.1573"